OpenStudy (anonymous):

wx^5 y^3-w^4 xy^4 complete factored form ?

OpenStudy (anonymous):

factor wxy^3

OpenStudy (anonymous):

\[wxy^3(x^4-w^3y)\]

OpenStudy (anonymous):

\[wx^5 y^3-w^4 xy^4=wxy^2(x^4-w^3y^2)=wxy^2(x^4-w^3y^2)\]

OpenStudy (anonymous):

is it wxy^2 or wxy^3 cause i think it is wxy^3

OpenStudy (anonymous):

@Jonask

OpenStudy (anonymous):

it is \[ wxy^2\] out side of the bracket.

OpenStudy (anonymous):

\[w^3 \] within the bracket

OpenStudy (anonymous):

w^3

OpenStudy (anonymous):

outside bracket \[wxy^3\]